OpenStudy (anonymous):

What is the degree measure of each interior angle of a regular triangle? 45° 60° 90° 120°

hartnn (hartnn):

total sum = 180 total angles = 3 so each measure will be?

hartnn (hartnn):

whats 180/3 ??

OpenStudy (mayankdevnani):

@janett77 what do you mean by this question?? Can you equate by a diagram...

OpenStudy (mayankdevnani):

@janett77 plz...response

OpenStudy (mayankdevnani):

@hartnn can you tell me

OpenStudy (anonymous):

@hartnnits 60

hartnn (hartnn):

yup, its 60.

OpenStudy (mayankdevnani):

plz,tell me @janett77

hartnn (hartnn):

mayank , regular triangle means equilateral triangle........

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

it's so simple...60 degree

OpenStudy (anonymous):

@mayankdevnani i really don't know what means thats why i asked but yes its 60

OpenStudy (mayankdevnani):

because we know that all sides of an regular triangle is equal.so angle opposite to equal sides are equal. so angles would be x+x+x=180 3x=180 x=180/3=60

OpenStudy (mayankdevnani):

@janett77

OpenStudy (anonymous):

i believe that is correct

OpenStudy (mayankdevnani):

yaa plz believe